OpenStudy (anonymous):

x^2+square root 19x + 15 over 4=0 completing the square

OpenStudy (tkhunny):

\(x^{2} + \sqrt{19}x + 15/4 = 0\) \(x^{2} + \sqrt{19}x\;+\;?? = -15/4\;+\;??\) What's stopping you from finding half of \(\sqrt{19}\) and then squaring it? It is the same as any other number in that position.

OpenStudy (anonymous):

what is half of \(\sqrt{19}\) ?